# Unifying Lane-Level Traffic Prediction from a Graph Structural Perspective: Benchmark and Baseline

### v1

@@ V-A1 PeMS Dataset
The PeMS dataset is a renowned open dataset system in the field of transportation, with its subsets PeMSD7 and PeMSD8 widely used as benchmarks in road-level traffic prediction. This system integrates a wealth of traffic information provided by the California Department of Transportation and other local agencies, covering data from traffic detectors, incident records, lane closures, toll tags, census traffic counts, and road inventories. Given the extensive information sources and reliability of the data in PeMS, we carefully selected 8 sensors that have been consistently operational with high data integrity from approximately 40,000 available sensors. These chosen sensors exhibit minimal data missingness, ensuring the quality and usability of the dataset. We used the speed and flow data recorded by these sensors from February 5 to March 5, 2017, as our baseline dataset for lane-level traffic prediction. For the few instances of missing data, we employed imputation using the mean values of adjacent time slots to maintain the continuity and completeness of the data. As illustrated in Figure 6 (c), the processed sensor data clearly displays significant daily cyclical variations, providing vital foundational information for analyzing and predicting lane-level traffic flow. Furthermore, to support the research needs of other scholars, we point out that researchers interested in a more in-depth study can directly access the PeMS system to obtain additional sensor data for the corresponding roads. This access not only facilitates broader research in the field but also encourages the use of reliable and comprehensive data in traffic studies.
@@ V-A2 HuaNan Dataset
To ensure the quality and reliability of the dataset, we conducted thorough cleaning and preprocessing of the extracted data. This included identifying and correcting anomalies, filling in missing values, and performing necessary data normalization procedures. This process ensured the completeness and consistency of the dataset, laying a solid foundation for subsequent analysis and modeling. In Figure 6, we present a trend analysis of all sensors data. These trends reflect the traffic flow characteristics of the HuaNan Expressway during different periods, providing an intuitive perspective on the usage patterns of the road and the characteristics of peak traffic periods. Additionally, the chosen road segment, devoid of entrances and exits, makes this dataset particularly suitable as a benchmark for prediction models that are only applicable to roads with a regular number of lanes.
@@ V-C Baselines and Code Configuration
We conducted a comparative analysis of fourteen models, including the GraphMLP model, designed for lane-level traffic prediction, and nine graph-structured spatio-temporal prediction models across three datasets. Given the scarcity of publicly available code for models proposed in the domain of lane-level traffic prediction, this section not only introduces the baseline models but also briefly describes the process of experimental code replication or the sources of the code.
@@ V-C1 Lane-level Traffic Prediction Model
  • HGCN[23] and DGCN[24]: The two models in question, originating from the same author, utilize identical formulas and methodologies, integrates spatial dependency analysis, data fusion, and temporal attention. Implemented using PyTorch, excluding heterogeneous data for fairness.
@@ V-C2 Graph-structured Spatio-temporal Prediction Models
TABLE I: Comparisons on PeMS dataset
Horizon 3 6 12
Metric MAE RMSE MAPE Cost MAE RMSE MAPE Cost MAE RMSE MAPE Cost
Cat-RF-LSTM 7.75 10.68 81.50% -- 7.81 11.00 81.92% -- 8.30 11.90 83.47% --
@@ V-E Lane-Level Traffic Benchmark
In Table I, we present the experimental results on the PeMS dataset. From the average values, we observe that the predictive performance of graph-structured models generally surpasses the average of lane-level models, indicating the feasibility of transferring graph-structured models to lane-level traffic prediction. However, the average training time for lane-level traffic prediction models is less than that for graph-structured models. This is attributed to the typically simpler structure of lane-level models, designed to adapt to the highly dynamic nature of lane traffic, presenting challenges for the direct transfer of graph-structured models to lane-level prediction. Among lane-level prediction models, MDL and ST-AFN, which utilized the authors’ code and experimental configurations, performed relatively well. Unexpectedly, the FDL model performed worse than LSTM and GRU, most likely due to the ineffective computation of its entropy-based grey relational algorithm. For graph-structured models, STGODE, STSGCN, GraphWave Net, and AGCRN all demonstrated commendable performance. Notably, STSGCN performed slightly better in the MAPE metric, indicating its effectiveness in predicting smaller actual values. STGCN was the fastest training model among graph-structured models, benefiting from its fully convolutional architecture. Overall, the GraphMLP model exhibited the best performance, surpassing the most effective STGODE model among all baselines in terms of prediction accuracy. Although its training time was not as short as some lane-level models, its higher predictive accuracy demonstrated its superior predictive effectiveness and efficiency.
TABLE II: Comparisons on HuaNan dataset
Horizon 3 6 12
Metric MAE RMSE MAPE Cost MAE RMSE MAPE Cost MAE RMSE MAPE Cost
Cat-RF-LSTM 18.17 22.71 93.50% -- 18.30 22.89 93.67% -- 18.67 23.21 99.23% --
@@
MDL $\underline{3.63}$ $\underline{5.86}$ $\underline{18.74\%}$ 5.22 $\underline{5.86}$ $\underline{9.01}$ $\underline{29.63\%}$ 5.31 $\underline{8.37}$ $\underline{12.56}$ $\underline{31.79\%}$ 5.77
@@
STSGCN 10.85 15.28 37.17% 8.56 11.76 16.28 40.49% 8.75 12.47 17.05 44.70% 8.95
AGCRN $\underline{4.11}$ $\underline{6.08}$ $\underline{16.72\%}$ 10.47 $\underline{5.28}$ $\underline{7.49}$ $\underline{21.70\%}$ 12.04 $\underline{6.47}$ $\underline{10.75}$ $\underline{24.81\%}$ 13.22
@@
Mean 11.71 16.23 42.50% 8.43 12.65 17.46 47.99% 9.17 13.74 18.93 51.22% 10.11
GraphMLP 2.99 4.03 11.53% 2.97 3.88 7.77 17.03% 3.09 4.88 9.00 21.20% 3.28
Table II demonstrates the performance of various models on the HuaNan dataset. Due to unit differences, the actual values in the HuaNan dataset are higher than those in the PeMS dataset, resulting in a significant amplification of errors. Observing the average performance, similar to the PeMS dataset, spatio-temporal graph-structured models outperform lane-level prediction models in terms of average prediction accuracy, yet require longer training times. The magnified errors highlight the distinctions between models more conspicuously. Among lane-level prediction models, MDL and ST-AFN remain the top performers, with MDL showing particularly notable performance. In contrast to the PeMS dataset, spatio-temporal graph-structured models exhibited some variances, where the AGCRN model surpassed STGODE, STSGCN, and GraphWave Net models across all three prediction horizons and metrics, emerging as the top-performing model. The MTGNN model also demonstrated commendable performance, while STGCN continued to have the least training time consumption. In a comprehensive assessment, GraphMLP maintained superior performance over all other baselines, exceeding the prediction accuracy of both the MDL and AGCRN models while requiring less training time. Moreover, GraphMLP showed more stable long-term predictive performance at horizons of 6 and 12, with its advantage in training time efficiency becoming even more pronounced.
TABLE III: Comparisons on PeMSF dataset
@@ V-F Comparison of Irregular Lane Performance
Table III showcases the performance of models that possess this capability. From the Difference index, we observe that most lane-level models exhibit stronger adaptability to scenarios with multiple lanes, with an average Difference index of -0.45%, indicating an improvement in performance on the irregular lane dataset compared to regular lane numbers. The average Difference index for spatio-temporal graph-structured models is 0.42%, suggesting that the average MAE on the irregular lane dataset, PeMSF, is larger compared to regular roads, indicating that some spatio-temporal graph models lack sufficient adaptability to irregular lane configurations. However, GraphMLP achieved a 3.93% reduction in MAE, surpassing the lane-level model STA-ED, which reduced by 2.28%, and the spatio-temporal graph-structured model MTGNN, which reduced by 2.32%, making it the best-performing model. This improvement is likely due to the additional vehicles providing more valuable adjacency information to other lane segments within the dynamic graph network of the GraphMLP model.
@@ V-G Performance on Long-Term Prediction
(a) Comparison of MAE on PeMS dataset
with Horizon=3
(b) Comparison of RMSE on PeMS dataset
with Horizon=3
(c) Comparison of MAE on HuaNan dataset
with Horizon=3
(d) Comparison of RMSE on HuaNan dataset
with Horizon=3
(e) Comparison of MAE on PeMS dataset
with Horizon=6
(f) Comparison of RMSE on PeMS dataset
with Horizon=6
(g) Comparison of MAE on HuaNan dataset
with Horizon=6
(h) Comparison of RMSE on HuaNan dataset
with Horizon=6
(i) Comparison of MAE on PeMS dataset
with Horizon=12
(j) Comparison of RMSE on PeMS dataset
with Horizon=12
(k) Comparison of MAE on HuaNan dataset
with Horizon=12
(l) Comparison of RMSE on HuaNan dataset
with Horizon=12
Figure 10: Visual analysis of model performance based on training cost and prediction accuracy
@@ V-I Ablation Study
Although our primary contribution does not lie in proposing a new model for lane-level traffic prediction, we deem it necessary to conduct a series of ablation studies to validate the impact of each component of the GraphMLP model on prediction accuracy and time consumption. Table IV presents the comparative results of the Instance Normalization module, the Dynamic Graph Network module, the Independent Temporal MLP module, and the complete model on both the PeMS and HuaNan datasets.
@@
TABLE IV: Ablation Study
@@
VI Conclusion
@@
In this paper, we conduct a detailed analysis of the unique challenges faced by lane-level traffic prediction compared to coarser-grained city and road-level traffic prediction, and provide a comprehensive review and categorization of current research in the field. Based on this analysis, we define a unified spatial topology and prediction task for lane-level traffic prediction and introduce a concise and efficient baseline model, GraphMLP. This model employs adaptively generated dynamic graph convolution and patch-based independent temporal MLP networks for efficient prediction. For further empirical analysis, we offer three public datasets, from two real road networks with both regular and irregular numbers of lanes. We have also replicated previously unpublished model codes and employed comprehensive evaluation metrics for a fair comparison of the accuracy and efficiency of existing models. Through this work, we aim to further advance the field of lane-level traffic prediction, providing richer perspectives and tools for traffic forecasting tasks.
Through the analysis presented in this paper, we posit that future research in lane-level traffic prediction should focus on enhancing prediction accuracy, reducing training times, and refining models for ease of fine-tuning. Accurate lane-level predictions can not only offer crucial support for more granular route planning, thereby aiding autonomous vehicles in devising quicker and safer travel paths, but can also provide more effective lane-changing recommendations for human-driven vehicles. Moreover, lane-level traffic prediction plays a significant role in advancing the development and refinement of large-scale models and pre-trained models within the traffic domain. Therefore, future studies should not only concentrate on technological advancements and improvements in model performance but also on how these advancements can be practically applied to enhance the overall efficiency and safety of transportation systems.

### v2

@@ V-A1 PeMS Dataset
The PeMS dataset is a renowned open traffic data system managed by the California Department of Transportation, widely used in road-level traffic studies. To construct a high-quality lane-level dataset, we carefully selected 8 sensors from approximately 40,000 available detectors, based on data continuity and operational reliability. The resulting dataset covers the period from February 5 to March 5, 2017, including lane-level speed and flow data. To address rare cases of missing values, we applied mean imputation using adjacent time slots. The data shows clear daily periodicity, as visualized in Figure 6(a), making it suitable for temporal traffic modeling. Researchers can also access the original PeMS platform for further exploration and extension.
V-A2 PeMSF Dataset
The PeMSF dataset is an extended variant of the PeMS dataset that includes freeway segments with entrance ramps, where the number of lanes increases from 5 to 6 at certain locations. This variation reflects more realistic and irregular lane-level structures common in urban expressways, especially near merges or exits. PeMSF is designed to test the adaptability of prediction models to dynamic and non-uniform lane configurations, an essential capability for deployment in complex real-world traffic networks. Figure 5(a) highlights these irregular segments.
@@ V-A3 HuaNan Dataset
To ensure the quality and reliability of the dataset, we thoroughly cleaned and preprocessed the extracted data. This included identifying and correcting anomalies, filling in missing values, and performing necessary data normalization procedures. This process ensured the completeness and consistency of the dataset, laying a solid foundation for subsequent analysis and modeling. In Figure 6, we present a trend analysis of all sensors data. These trends reflect the traffic flow characteristics of the HuaNan Expressway during different periods, providing an intuitive perspective on the usage patterns of the road and the characteristics of peak traffic periods. Additionally, the chosen road segment, devoid of entrances and exits, makes this dataset particularly suitable as a benchmark for prediction models that are only applicable to roads with a regular number of lanes.
@@ V-A4 Statistical Analysis
To visually present and compare the core characteristics of our publicly available datasets, we summarize the detailed information of the three datasets in Table III, providing a systematic comparative analysis.
TABLE III: Statistics of Datasets
Parameters PeMS PeMSF HuaNan
Timespan 2/5/2017- 2/5/2017- 7/22/2022-
3/5/2017 3/5/2017 8/22/2022
Region Los Angeles Los Angeles Guangzhou
# of Objects 322,360 346,537 3,214,080
# of Static Edges 67 70 121
Road Type Freeway Freeway Urban Expressway
Unit Miles/Hour Miles/Hour Kilometers /Hour
Time Interval 5 min 5 min 2min
@@ V-C Baselines and Code Configuration
We conducted a comparative analysis of fourteen models, including the GraphMLP model, designed for lane-level traffic prediction, and nine graph-structured spatio-temporal prediction models across three datasets. Given the general lack of publicly available code for models proposed in the field of lane-level traffic prediction, this section not only introduces the baseline model but also briefly outlines the process of reproducing the experimental code and its source. The input and output of the reproduction framework remain consistent with those in Algorithm 1.
@@ V-C1 Lane-level Traffic Prediction Model
  • HGCN[26] and DGCN[27]: The two models in question, originating from the same author, utilize identical formulas and methodologies, integrate spatial dependency analysis, data fusion, and temporal attention. Implemented using PyTorch, excluding heterogeneous data for fairness.
@@ V-C2 Graph-structured Spatio-temporal Prediction Models
TABLE IV: Comparisons on PeMS dataset
Model Horizon 3 6 12
Type Metric MAE RMSE MAPE Cost MAE RMSE MAPE Cost MAE RMSE MAPE Cost
Lane-Level Cat-RF-LSTM 7.75 10.68 81.50% -- 7.81 11.00 81.92% -- 8.30 11.90 83.47% --
@@ V-E Lane-Level Traffic Benchmark
In Table IV, we present the experimental results on the freeway dataset PeMS. From the average values, we observe that the predictive performance of graph-structured models generally surpasses the average of lane-level models, indicating the feasibility of transferring graph-structured models to lane-level traffic prediction. However, the average training time for lane-level traffic prediction models is less than that for graph-structured models. This is attributed to the typically simpler structure of lane-level models, designed to adapt to the highly dynamic nature of lane traffic, presenting challenges for the direct transfer of graph-structured models to lane-level prediction. Among lane-level prediction models, MDL and ST-AFN, which utilized the authors’ code and experimental configurations, performed relatively well. Unexpectedly, the FDL model performed worse than LSTM and GRU, most likely due to the ineffective computation of its entropy-based grey relational algorithm. For graph-structured models, STGODE, STSGCN, GraphWaveNet, and AGCRN all demonstrated commendable performance. Notably, STSGCN performed slightly better in the MAPE metric, indicating its effectiveness in predicting smaller actual values. STGCN was the fastest training model among graph-structured models, benefiting from its fully convolutional architecture. Overall, the GraphMLP model exhibited the best performance, surpassing the most effective STGODE model among all baselines in terms of prediction accuracy. Although its training time was not as short as some lane-level models, its higher predictive accuracy demonstrated its superior predictive effectiveness and efficiency.
TABLE V: Comparisons on HuaNan dataset
Model Horizon 3 6 12
Type Metric MAE RMSE MAPE Cost MAE RMSE MAPE Cost MAE RMSE MAPE Cost
Lane-Level Cat-RF-LSTM 18.17 22.71 93.50% -- 18.30 22.89 93.67% -- 18.67 23.21 99.23% --
@@
MDL $\underline{3.63}$ $\underline{5.86}$ $\underline{18.74\%}$ 5.22 $\underline{5.86}$ $\underline{9.01}$ $\underline{29.63\%}$ 5.31 $\underline{8.37}$ $\underline{12.55}$ $\underline{31.79\%}$ 5.77
@@
STSGCN 10.85 15.28 37.17% 8.56 11.76 16.28 40.49% 8.75 12.47 17.05 44.70% 89.53
AGCRN $\underline{4.11}$ $\underline{6.08}$ $\underline{16.72\%}$ 10.47 $\underline{5.28}$ $\underline{7.49}$ $\underline{21.70\%}$ 24.04 $\underline{6.47}$ $\underline{10.75}$ $\underline{24.81\%}$ 13.22
@@
Mean 11.71 16.23 42.50% 8.43 12.65 17.46 47.99% 12.17 13.74 18.93 51.22% 20.19
New GraphMLP 2.99 4.03 11.53% 2.97 3.88 7.77 17.03% 3.09 4.88 9.00 21.20% 3.28
Table V demonstrates the performance of various models on the urban expressway HuaNan dataset. Due to unit differences, the actual values in the HuaNan dataset are higher than those in the PeMS dataset, resulting in a significant amplification of errors. Observing the average performance, similar to the PeMS dataset, spatio-temporal graph-structured models outperform lane-level prediction models in terms of average prediction accuracy, yet require longer training times. The magnified errors highlight the distinctions between models more conspicuously. Among lane-level prediction models, MDL and ST-AFN remain the top performers, with MDL showing particularly notable performance. In contrast to the PeMS dataset, spatio-temporal graph-structured models exhibited some variances, where the AGCRN model surpassed STGODE, STSGCN, and GraphWave Net models across all three prediction horizons and metrics, emerging as the top-performing model. The MTGNN model also demonstrated commendable performance, while STGCN continued to have the least training time consumption. In a comprehensive assessment, GraphMLP maintained superior performance over all other baselines, exceeding the prediction accuracy of both the MDL and AGCRN models while requiring less training time. Moreover, GraphMLP showed more stable long-term predictive performance at horizons of 6 and 12, with its advantage in training time efficiency becoming even more pronounced. Compared to the results on the PeMS dataset, the higher computational cost reflects the greater complexity of urban roads, which demands more refined spatio-temporal modeling capabilities from the models.
TABLE VI: Comparisons on PeMSF dataset
@@ V-F Comparison of Irregular Lane Performance
Table VI showcases the performance of models that possess this capability. From the Difference index, we observe that most lane-level models exhibit stronger adaptability to scenarios with multiple lanes, with an average Difference index of -0.45%, indicating an improvement in performance on the irregular lane dataset compared to regular lane numbers. The average Difference index for spatio-temporal graph-structured models is 0.42%, suggesting that the average MAE on the irregular lane dataset, PeMSF, is larger compared to regular roads, indicating that some spatio-temporal graph models lack sufficient adaptability to irregular lane configurations. However, GraphMLP achieved a 3.93% reduction in MAE, surpassing the lane-level model STA-ED, which reduced by 2.28%, and the spatio-temporal graph-structured model MTGNN, which reduced by 2.32%, making it the best-performing model. This improvement is likely due to the additional vehicles providing more valuable adjacency information to other lane segments within the dynamic graph network of the GraphMLP model.
@@ V-I Ablation Study
Although our primary contribution does not lie in proposing a new model for lane-level traffic prediction, we deem it necessary to conduct a series of ablation studies to validate the impact of each component of the GraphMLP model on prediction accuracy and time consumption. Table VII presents the comparative results of the Instance Normalization module, the Dynamic Graph Network module, the Independent Temporal MLP module, and the complete model on both the PeMS and HuaNan datasets.
@@
TABLE VII: Ablation Study
@@
V-J Case Study
To further demonstrate the characteristics of lane-level traffic prediction and intuitively evaluate the performance of the simple baseline model, GraphMLP in complex scenarios, we conduct a case study on a representative segment from the PeMSF dataset, which features irregular lane configurations. Specifically, we select the sixth lane segment $l_{4,6}$ under the fourth sensor and visualize its topological position and prediction results under a historical window of $T=12$ and a prediction horizon of $z=12$, as shown in Figure 10.
Subfigure (a) shows the topology around $l_{4,6}$, where entrance lanes and other asymmetric structures are clearly present. These irregular configurations are difficult to capture using grid-based models, whereas the graph-based approach flexibly represents lane connectivity, including dynamic structures such as entrance ramps and construction lanes, offering a more realistic and expressive modeling capability. Subfigure (b) compares the historical, actual future, and predicted speeds of the selected lane segment. GraphMLP effectively tracks traffic dynamics, with predictions closely matching the ground truth. The narrow error band suggests strong fitting ability and good generalization. This case highlights the practicality of graph-based representations for modeling irregular lane structures and demonstrates the adaptability of GraphMLP, as a baseline model, across diverse traffic scenarios.
(a) Topological position of lane segment $l_{4,6}$
(b) Prediction error of lane segment $l_{4,6}$
Figure 10: Case study on the topological structure and prediction visualization of lane-level traffic prediction on irregular roads.
(b) Comparison of MAE on PeMS dataset
with Horizon=3
(c) Comparison of RMSE on PeMS dataset
with Horizon=3
(d) Comparison of MAE on HuaNan dataset
with Horizon=3
(e) Comparison of RMSE on HuaNan dataset
with Horizon=3
(f) Comparison of MAE on PeMS dataset
with Horizon=6
(g) Comparison of RMSE on PeMS dataset
with Horizon=6
(h) Comparison of MAE on HuaNan dataset
with Horizon=6
(i) Comparison of RMSE on HuaNan dataset
with Horizon=6
(j) Comparison of MAE on PeMS dataset
with Horizon=12
(k) Comparison of RMSE on PeMS dataset
with Horizon=12
(l) Comparison of MAE on HuaNan dataset
with Horizon=12
(m) Comparison of RMSE on HuaNan dataset
with Horizon=12
Figure 11: Visual analysis of model performance based on training cost and prediction accuracy. The red and blue lines represent the mean values of the respective metrics.
VI Conclusion and Future Work Discussion
@@
In this paper, we conduct a detailed analysis of the unique challenges faced by lane-level traffic prediction compared to coarser-grained city and road-level traffic prediction and provide a comprehensive review and categorization of current research in the field. Based on this analysis, we define a unified spatial topology and prediction task for lane-level traffic prediction and introduce a concise and efficient baseline model, GraphMLP. This model employs adaptively generated dynamic graph convolution and patch-based independent temporal MLP networks for efficient prediction. For further empirical analysis, we offer three public datasets from two real road networks with both regular and irregular numbers of lanes. We have also replicated previously unpublished model codes and employed comprehensive evaluation metrics for a fair comparison of the accuracy and efficiency of existing models. Through this work, we aim to further advance the field of lane-level traffic prediction, providing richer perspectives and tools for traffic prediction tasks.
From the analysis presented in this paper, we believe current research still faces several limitations, such as the difficulty in modeling fine-grained and rapidly changing lane-level dynamics, limited generalizability across diverse road types, and the high computational cost of real-time inference. Therefore, future research in lane-level traffic prediction should focus on the following key areas to drive technological advancement and practical applications:
Efficiency Design: Future research should prioritize improving the efficiency of lane-level traffic prediction models, particularly in terms of training time and computational resource consumption. By optimizing model architectures, adopting more efficient training methods, or utilizing model compression techniques, the computational complexity of the models can be reduced, thereby enhancing their response speed and processing capacity in real-world applications. This would not only improve the performance of real-time traffic prediction but also enable lane-level prediction to be applied more broadly in intelligent transportation systems.
Road-Lane Interactions: As the complexity of traffic networks increases, future research needs to explore the interactions between different lanes in greater depth. In real-world traffic scenarios, changes in flow, traffic mobility, and lane-switching behaviors are closely interrelated. Therefore, effectively modeling these interactions at the lane level and integrating spatiotemporal information is key to improving prediction accuracy. Meanwhile, such modeling also opens opportunities for discovering latent traffic structures and bottlenecks through interpretable data mining. Future studies can explore closer road-lane interactions to further capture and model the complex dependencies between lanes.
Multi-Granularity Collaboration: Future lane-level traffic prediction should not only focus on predictions at a single granularity level but also consider the collaborative functioning of multiple granularities (e.g., road-level, region-level, and city-level). Multi-granularity models can leverage traffic information at different levels to achieve more comprehensive and accurate predictions. Especially in complex urban environments, the coordination between coarse-grained road-level predictions and fine-grained lane-level predictions will significantly enhance overall prediction performance.
Large-Scale Traffic Models: Research on lane-level traffic prediction needs to better integrate into the development of large-scale traffic models, particularly in the context of analyzing large-scale road networks and long-duration datasets. As data volumes grow and model complexities increase, efficiently training and optimizing these large models will be an important direction for future research. Techniques such as pre-trained models and transfer learning can improve model generalization, reduce dependence on vast amounts of labeled data, and enhance the predictive capabilities of large-scale intelligent transportation systems.
Lane-level traffic prediction is a more complex and detailed task. Future research should focus not only on model innovation and performance optimization but also on practical deployment in real-world traffic systems to improve efficiency, safety, and adaptabilit, such as providing lane-level guidance for autonomous vehicles, supporting dynamic lane control policies like reversible (tidal) lanes, enabling proactive congestion mitigation at fine granularity, and assisting urban traffic signal optimization. Integrating external factors like traffic signals and lane-changing behavior will be crucial for advancing intelligent traffic management, enabling more accurate predictions, optimizing traffic flow, and supporting autonomous driving systems.